# X-SAM: From Segment Anything to Any Segmentation

### v1

@@ Abstract
1 Introduction
Multi-modal Large Language Models (MLLMs) have exhibited substantial advancements alongside the rapid development of Large Language Models (LLMs) (Bai et al. 2023; Touvron et al. 2023a, b; Abdin et al. 2024) and multi-modal pre-training methods (Radford et al. 2021; Jia et al. 2021; Zhai et al. 2023). These models have shown remarkable effectiveness in a wide range of applications, including image captioning (Xu et al. 2015), VQA (Antol et al. 2015), and visual editing (Chen et al. 2018). Nevertheless, a significant impediment to developing a truly generalized model remains: contemporary MLLMs are restricted to generating solely textual outputs. This limitation poses a considerable challenge in directly addressing tasks that require pixel-level comprehension of visual data, such as image segmentation, which is the most critical task in the field of computer vision.
@@ 1 Introduction
Fig. 1: Illustration of the capabilities of X-SAM. (a). Text query tasks: Generic(Gen.), Referring(Ref.), Reasoning(Rea.), and Grounded Conversation Generation(GCG) segmentation, etc.. (b). Vision query tasks: Interactive(Inter.) and Visual GrounDed(VGD) segmentation for single and cross-image. (c). X-SAM outperforms existing MLLMs on all segmentation benchmarks. Best viewed in zoom.
@@
First, we develop a unified segmentation MLLM architecture that incorporates a unified mask decoder capable of generating segmentation masks suitable for generalized image segmentation tasks. Second, we expand the multimodal capabilities of MLLMs to process not only textual queries but also visual queries. Specifically, we introduce a novel task termed Visual Grounded (VGD) segmentation, which segments all instance objects with interactive visual prompts in an image. This task introduces visual guide modalities into large language models (LLMs). Furthermore, we propose a unified input format and training methodology that reformulates segmentation tasks within a unified framework, thus optimizing the adaptation of MLLMs to diverse image segmentation tasks.
As shown in Fig. 1 and Tab. 1, we present the comprehensive capabilities of X-SAM and compare them with those of other methods. Our proposed framework exhibits capabilities in processing text query-based tasks, such as generic segmentation and referring segmentation, while simultaneously accommodating vision query-based tasks such as interactive segmentation (Zhang et al. 2024d) and our novel VGD segmentation, which functions effectively in both single-image and cross-image contexts. Furthermore, X-SAM leverages the reasoning and generative capacities of LLMs, thereby enabling advanced reasoning segmentation and Grounded Conversation Generation (GCG) (Rasheed et al. 2024) segmentation.
X-SAM undergoes co-training with a diverse range of datasets. We perform a comprehensive evaluation on more than twenty segmentation datasets across seven distinct image segmentation tasks, even including the image conversion task. X-SAM achieves the state-of-the-art performance across all image segmentation benchmarks, and establishes a robust new baseline for unified pixel-level image understanding. In summary, our contributions are as follows:
@@
2 Related Work
Multi-modal Large Language Model. Multi-modal learning has evolved from early models focused on task-specific fusion and feature extraction (Li et al. 2022b), to leveraging large language models (Brown et al. 2020; Touvron et al. 2023a, b) for generalized, instruction-tuned multi-task benchmarks (Li et al. 2023a; Liu et al. 2024c; Hudson et al. 2019). LLaVA (Liu et al. 2024a, b, 2023a) introduced visual feature tokenization, inspiring advances in visual representation (Yuan et al. 2024b), specialized vision extensions (Lai et al. 2024; Lin et al. 2023; Dong et al. 2024a, b; Zhang et al. 2023; Ren et al. 2024; Zhang et al. 2024a; Zang et al. 2025), and language-guided segmentation (Li et al. 2024e; Zhang et al. 2024b). However, most progress remains task-specific. To our knowledge, we are the first to successfully implement a comprehensive approach, opening new directions for image segmentation.
Tab. 1: Comparison of Capability. We compare different methods on both segmentation-specific (Gray) and MLLM-based.
@@ 2 Related Work
PixelLM(Zhang et al. 2024d)
OMG-LLaVA(Zhang et al. 2024c)
@@
Multi-modal Grounded Segmentation. Recent works (Pan et al. 2024; Sun et al. 2024; Zhou et al. 2022a; Bar et al. 2022; Wang et al. 2023a, b) explore visual initiation methods in vision, including learnable tokens (Zhou et al. 2022a), mask-visual-modeling (Wang et al. 2024; Fang et al. 2023; Wang et al. 2023b), and visual prompting encoders (Yuan et al. 2024a; Wang et al. 2023c, a). SAM (Kirillov et al. 2023b) and its extensions (Xu et al. 2024; Yuan et al. 2024a) introduce visual grounding signals to segmentation models, greatly improving performance. Interactive segmentation (Li et al. 2024e) further enhances user-guided segmentation for MLLMs. However, existing methods cannot freely treat grounded input as textual input for segmentation. To address this, we propose Visual GrounDed (VGD) segmentation, enabling more diverse multi-modal grounded segmentation.
Unified Segmentation Model. Vision transformers (Mehta et al. 2021; Dosovitskiy et al. 2020; Carion et al. 2020) have advanced universal segmentation, with recent works (Zhou et al. 2022b; Xu et al. 2024; Li et al. 2024d; Zhou et al. 2024; Sun et al. 2023; Yang et al. 2021; Cheng et al. 2022b; Wang et al. 2021a) developing end-to-end mask classification frameworks that outperform earlier models (Zhou et al. 2022c; Li et al. 2021, 2020; Chen et al. 2019) across various applications. Research has expanded to open-world and open-vocabulary segmentation (Wu et al. 2024; Yuan et al. 2024a; Qi et al. 2022a, b), as well as unified architectures for multiple tasks (Athar et al. 2023; Gu et al. 2023; Yan et al. 2023b; Xu et al. 2024; Jain et al. 2023; Li et al. 2024e). However, most methods focus solely on visual segmentation and lack interactive textual and visual prompts found in MLLMs. To address this, we combine SAM with MLLMs, extending SAM from segment anything to any segmentation, and introduce a unified framework adaptable to all image segmentation tasks, establishing a new strong baseline.
3 Method
@@ 3.1 Formulation
Vision Query Input. Beyond text query inputs, some tasks necessitate vision query inputs, such as interactive segmentation (Zhang et al. 2024d) and the Visual GrounDed segmentation proposed in this work. In contrast to text query inputs, vision query inputs incorporate a visual prompt from the user, which may take the form of points, scribbles, boxes, or masks. To denote the visual prompt, we employ a dedicated token, <region>, within the input format. Analogous to the text query input, the visual prompt is formatted as “<p><region></p>” and the segmentation output is similarly indicated by the <SEG> token. The <region> token serves as a placeholder for the visual prompt and will be replaced by the region feature extracted from the segmentation encoder.
@@
Unified Formulation. The latent language embeddings between the <p> and </p> tokens are used as the condition embeddings for the segmentation decoder to compute the classification scores. Based on this formulation, we achieve a unified framework for all image segmentation tasks. Given an input image $\mathbf{X}_{\mathrm{v}}\in\mathbb{R}^{H\times W\times 3}$ and a language instruction $\mathbf{X}_{\mathrm{q}}\in\mathbb{R}^{P\times 1}$, the model takes the image and language instruction as inputs and outputs a language response $\mathbf{Y}_{\mathrm{q}}\in\mathbb{R}^{L\times 1}$ and a segmentation mask $\mathbf{Y}_{\mathrm{m}}\in\mathbb{R}^{H\times W}$. Here, $P$ is the length of the input text tokens, and $L$ is the total length of the input and output text tokens. $H$ and $W$ denote the height and width of the image, respectively. Detailed input format examples can be found in Fig. 1 (a) and (b).
Fig. 2: The Overview of X-SAM. X-SAM comprises dual encoders, dual projectors, a language model, a segmentation connector, and a segmentation decoder. The dual encoders process the image and project features to match text embedding dimensions, which are then input to the language model with tokenized text for instruction-guided understanding. The SAM features are connected to the segmentation decoder, which uses the LLM’s <SEG> token to generate segmentation masks.
3.2 Architecture
@@
In this section, we propose X-SAM, a unified segmentation MLLM for any segmentation. As shown in Fig. 2, it includes dual encoders, dual projectors, an LLM, a segmentation connector, and a segmentation decoder.
Dual Projectors. To enhance the LLM’s understanding of the image, we concatenate the features from the image encoder and the segmentation encoder before passing them to the LLM. Specifically, the feature from the segmentation encoder is too large to be processed directly by the LLM, so we utilize a pixel-shuffle operation to reduce its spatial size. We then project the reduced feature into the language embedding space $\mathbf{H}_{\textit{q}}$ via an MLP projector $\mathbf{W}_{\textit{s}}$. For the feature from the image encoder, we directly project it into the language embedding space via an MLP projector $\mathbf{W}_{\textit{i}}$, such that $\mathbf{H}_{\textit{v}}=\mathbf{W}_{\textit{i}}\cdot\mathbf{Z}_{\texttt{v}}$ and $\mathbf{H}_{\textit{s}}=\mathbf{W}_{\textit{s}}\cdot\mathbf{Z}_{\texttt{s}}$. We then concatenate the features from dual projectors and the language embeddings, and input them into the LLM $f_{\phi}$.
Fig. 3: The Architecture of Segmentation Connector.
Segmentation Connector. For image segmentation tasks, fine-grained multi-scale features are crucial for the segmentation decoder to accurately predict segmentation masks. The output of the segmentation encoder in SAM is single-scale (1/16) with reduced spatial resolution. To obtain multi-scale features, we design a segmentation connector $g_{\texttt{c}}$, to bridge the segmentation encoder and decoder. As shown in Fig. 3, we perform patch-merge using a pixel-shuffle (Chen et al. 2024) with a scale of 0.5 to reduce the spatial size of the last feature in the encoder to a smaller scale (1/32). We also perform patch-expand with a pixel-shuffle of scale 2.0 to increase the spatial size of the last feature to a larger scale (1/8), resulting in multi-scale features for the segmentation decoder.
Tab. 2: Comprehensive Performance Comparison. We compare X-SAM to segmentation-specific models (Gray) and MLLMs. “✗” denotes unsupported tasks. “–” indicates unreported results. X-SAM achieves state-of-the-art performance across all segmentation tasks with a single model. Best results are in bold, second-best are underlined.
@@
Segmentation Decoder. The Segment Anything Model (SAM) can segment a single object based on input text or visual prompts, but it fails to segment all objects in a single inference. To segment all objects at once, we replace its original segmentation decoder with a new decoder, following the approach in (Cheng et al. 2022a; VS et al. 2024). The segmentation decoder $g_{\psi}$ predicts masks and their category probabilities from either the input latent embedding $\mathbf{E}_{\mathrm{i}}$ or the output latent embedding $\mathbf{E}_{\texttt{o}}$, multi-scale segmentation features $\mathbf{F}_{\mathrm{c}}$, and a set of mask query tokens plus the <SEG> token embedding, which bridges the LLM output with the segmentation decoder. Notably, we introduce a latent background embedding to represent the “ignore” category for all tasks, thereby unifying all image segmentation tasks with one model.
3.3 Training
@@
Stage 2: Alignment Pre-training. To align the language embeddings and visual embeddings, we perform alignment pre-training on the LLaVA-558K dataset, following (Liu et al. 2023b). We keep the dual encoders and the LLM parameters frozen, and only train the dual projectors. In this way, the image embeddings and segmentation embeddings can be aligned with the pre-trained LLM word embeddings. The training objective for alignment pre-training is an auto-regressive loss $\mathcal{L}_{\mathrm{regressive}}$:
@@
4 Experiments
Tab. 3: Comparison of Referring Segmentation. We evaluate methods on referring segmentation benchmarks by (M)LLMs.
Method (M)LLM RefCOCO RefCOCO+ RefCOCOg
val testA testB val testA testB val test
SEEM-L (Zou et al. 2023b) - - - - - - - 65.6 -
UNINEXT-L (Yan et al. 2023a) - 80.3 82.6 77.8 70.0 74.9 62.6 73.4 73.7
UNINEXT-H (Yan et al. 2023a) - 82.2 83.4 81.3 72.5 76.4 66.2 74.7 76.4
OMG-LLaVA (Zhang et al. 2024c) InternLM-7B 77.2 79.8 74.1 68.7 73.0 61.6 71.7 71.9
Sa2VA(Yuan et al. 2025) InternVL2-8B 81.6 - - 76.2 - - 78.7 -
PSALM (Zhang et al. 2024d) Phi-1.5-1.3B 83.6 84.7 81.6 72.9 75.5 70.1 73.8 74.4
X-SAM (Ours) Phi-3-3.8B 85.1 87.1 83.4 78.0 81.0 74.4 83.8 83.9
Tab. 4: Comparison of GCG Segmentation. $\dagger$ indicates pretraining with the GranD dataset (Rasheed et al. 2024).
@@ 4 Experiments
4.1 Experiment Settings
Datasets and Tasks. For segmentor fine-tuning, we train on the COCO-Panoptic (Kirillov et al. 2019) dataset. For alignment pre-training, we utilize the LLaVA-558K (Liu et al. 2023b) dataset. For end-to-end mixed fine-tuning, we incorporate one image conversation dataset and five types of image segmentation datasets into the training process. To balance the training data across these diverse datasets, we set the training epoch to 1 and adjust the resampling rates of different datasets using dataset balance resampling. After training, X-SAM is capable of performing a variety of tasks, including Image Conversation, Generic, Referring, Reasoning, GCG, Interactive, and VGD Segmentation. Additionally, X-SAM supports Open-Vocabulary (OV) (OV-semantic, OV-instance, OV-panoptic) segmentation, enabling it to segment all objects defined by the input prompt, even those never seen before. Note that COCO-VGD is our proposed VGD segmentation dataset, which is built on the COCO2017 dataset. Details of the datasets are presented in Appendix A.1.
Evaluation Metrics. We conduct extensive experiments to evaluate the performance of X-SAM. For generic segmentation and open-vocabulary segmentation, we use PQ, mIoU, and mAP as the main metrics for panoptic, semantic, and instance segmentation, respectively. For referring segmentation and reasoning segmentation, we adopt cIoU and gIoU as metrics, following (Zhang et al. 2024d). For GCG segmentation, we use M, C, AP50, and mIoU as metrics, following (Rasheed et al. 2024). For interactive segmentation, we use mIoU and cIoU, also following (Zhang et al. 2024d). For VGD segmentation, we use AP and AP50. For image conversation, we adopt scores from common MLLM benchmarks as the main metrics, following (Liu et al. 2023b).
Implementation Details. We adopt the XTuner (Contributors 2023) codebase for training and evaluation. During segmentor fine-tuning, we train all parameters, set the batch size to 64, and use a learning rate of 1e-5 for the SAM encoder and 1e-4 for the other parameters. The number of training epochs is set to 36. For alignment pre-training, we train only the dual projector parameters, with a batch size of 256, a learning rate of 1e-3, and one training epoch. For end-to-end mixed fine-tuning, we train all parameters, set the batch size to 64, and use a learning rate of 4e-6 for the dual encoders and 4e-5 for the other parameters, with one training epoch. All training is conducted on 16 A100 GPUs. For image conversation evaluation, we use the VLMEvalKit (Duan et al. 2024) codebase to evaluate performance on MLLM benchmarks. For segmentation task evaluation, we follow the settings described in the corresponding papers and repositories. More implementation details are provided in Appendix A.3.
4.2 Main Results
Tab. 5: Comparison of VGD Segmentation. $\dagger$ indicates evaluation results following X-SAM setting.
@@ 4.2 Main Results
Tab. 6: Ablation on Fine-Tuning(FT).
@@
Tab. 7: Ablation on Dual Encoders. Swin is initialized from Mask2Former (M2F) (Cheng et al. 2022a).
@@
Overall. In Tab. 2, we compare X-SAM with current segmentation-specific models and MLLMs. X-SAM demonstrates the most comprehensive capabilities. It achieves performance comparable to state-of-the-art in generic segmentation, and achieves the best performance on other benchmarks, with a single model. X-SAM sets a new state-of-the-art record for image segmentation benchmarks. Detailed results for each task are discussed below.
Referring Segmentation. We evaluate X-SAM on RefCOCO, RefCOCO+, and RefCOCOg, with the results shown in Tab. 3. X-SAM outperforms PSALM (Zhang et al. 2024d) by 1.5% cIoU, 5.1% cIoU, and 10.0% cIoU on the validation sets of RefCOCO, RefCOCO+, and RefCOCOg, respectively. Compared to Sa2VA-8B (Yuan et al. 2025), X-SAM achieves better results with a smaller model size. It shows performance improvements of 3.5% cIoU, 1.8% cIoU, and 5.1% cIoU on RefCOCO, RefCOCO+, and RefCOCOg, respectively.
GCG Segmentation. Grounded conversation generation demands detailed image and pixel-level understanding, requiring MLLMs to link captioned objects to their segmentation masks. As shown in Tab. 4, X-SAM achieves a significant performance improvement compared to previous methods and obtains the best results on both the Val and Test sets. In terms of image-level understanding, X-SAM outperforms GLaMM (Rasheed et al. 2024) by 0.2% METEOR and 3.2% CIDEr on the Val set, and by 0.5% METEOR and 4.8% CIDEr on the Test set. In terms of pixel-level understanding, X-SAM outperforms OMG-LLaVA (Zhang et al. 2024c) by 3.3% AP and 3.9% mIoU on the Val set, and by 4.3% AP and 4.3% mIoU on the Test set.
VGD Segmentation. Visual grounded segmentation demands vision query understanding, requiring MLLMs to comprehend the visual modality and segment all related instances. Tab. 5 presents the VGD segmentation results. As VGD segmentation is our newly proposed task, we evaluate PSALM (Zhang et al. 2024d) following X-SAM’s settings. X-SAM outperforms PSALM by 45.9% AP, 45.9% AP, 45.8% AP, and 47.4% AP on Point, Scribble, Box, and Mask visual prompts, respectively.
More results and discussions for other segmentation and conversation benchmarks are provided in Appendix A.5.
4.3 Abaltions

### v2

@@ Abstract
Introduction
Figure 1: Illustration of Performance of X-SAM on Image Segmentation Benchmarks. X-SAM consistently surpasses existing Multimodal Large Language Models (MLLMs) across all evaluated segmentation benchmarks.
Multi-modal Large Language Models (MLLMs) have exhibited substantial advancements alongside the rapid development of Large Language Models (LLMs) (Bai et al. 2023; Touvron et al. 2023) and multi-modal pre-training methods (Radford et al. 2021; Jia et al. 2021). These models have shown remarkable effectiveness in a wide range of applications, including image captioning (Xu et al. 2015), VQA (Antol et al. 2015), and visual editing (Chen et al. 2018). However, current MLLMs lack the capability to generate dense pixel-level outputs for precise spatial understanding. This limitation poses a considerable challenge in directly addressing tasks that require pixel-level comprehension of visual data, such as image segmentation, which is the most critical task in the field of computer vision.
@@ Introduction
Figure 2: Illustration of the capabilities of X-SAM. (a). Text query tasks: Generic (Gen.), Referring(Ref.), Reasoning(Rea.), and Grounded Conversation Generation(GCG) segmentation, etc.. (b). Vision query tasks: Interactive(Inter.) and Visual GrounDed (VGD) segmentation for single and cross-image.
@@
First, we develop a unified segmentation MLLM architecture that incorporates a unified mask decoder capable of generating segmentation masks suitable for generalized image segmentation tasks. Second, we expand the multimodal capabilities of MLLMs to process not only textual queries but also visual queries. Specifically, we introduce a novel task termed Visual GrounDed (VGD) segmentation, which segments all instance objects with interactive visual prompts in an image. This task introduces visual guide modalities into large language models (LLMs). Moreover, we propose a unified input format and training methodology that reformulates segmentation tasks within a unified framework, thus optimizing the adaptation of MLLMs to diverse image segmentation tasks.
As shown in Figure 2 and Table 1, we present the comprehensive capabilities of X-SAM and compare them with those of other methods. Our proposed framework exhibits capabilities in processing text query-based tasks, such as generic segmentation and referring segmentation, while simultaneously accommodating vision query-based tasks such as interactive segmentation (Zhang et al. 2024c) and our novel VGD segmentation, which functions effectively in both single-image and cross-image contexts. Furthermore, X-SAM leverages the reasoning and generative capacities of LLMs, thereby enabling advanced reasoning segmentation and Grounded Conversation Generation (GCG) (Rasheed et al. 2024) segmentation.
X-SAM undergoes co-training with a diverse range of datasets. We perform a comprehensive evaluation on more than twenty segmentation datasets across seven distinct image segmentation tasks, even including the image conversion task. X-SAM achieves the state-of-the-art performance across all image segmentation benchmarks, and establishes a robust new baseline for unified pixel-level image understanding, as illustrated in Figure 1. In summary, our contributions are as follows:
@@
Related Work
@@
PixelLM(Zhang et al. 2024c)
OMG-LLaVA(Zhang et al. 2024b)
@@
Table 1: Comparison of Capability. We compare different methods on both segmentation-specific (Gray) and MLLM-based.
Multi-modal Large Language Model. Multi-modal learning has evolved from early models focused on task-specific fusion and feature extraction (Li et al. 2022b), to leveraging large language models (Brown et al. 2020; Touvron et al. 2023) for generalized, instruction-tuned multi-task benchmarks (Liu et al. 2024b; Hudson and Manning 2019). LLaVA (Liu et al. 2023b, 2024a) introduced visual feature tokenization, inspiring advances in visual representation (Yuan et al. 2024b), specialized vision extensions (Lai et al. 2024; Lin et al. 2023), and language-guided segmentation (Li et al. 2024d; Zhang et al. 2024a). However, most progress remains task-specific. To our knowledge, we are the first to successfully implement a comprehensive approach, opening new directions for image segmentation.
Multi-modal Grounded Segmentation. Recent works (Pan et al. 2024; Wang et al. 2023) explore visual initiation methods in vision, including learnable tokens (Zhou et al. 2022a), mask-visual-modeling (Fang et al. 2023), and visual prompting encoders (Yuan et al. 2024a). SAM (Kirillov et al. 2023b) and its extensions (Xu et al. 2024) introduce visual grounding signals to segmentation models, greatly improving performance. Interactive segmentation (Li et al. 2024d) further enhances user-guided segmentation for MLLMs. However, existing methods cannot freely treat grounded input as textual input for segmentation. To address this, we propose Visual GrounDed (VGD) segmentation, enabling more diverse multi-modal grounded segmentation.
Unified Segmentation Model. Vision transformers (Dosovitskiy et al. 2020) have advanced universal segmentation, with recent works (Xu et al. 2024; Cheng et al. 2022b) developing end-to-end mask classification frameworks that outperform earlier models (Zhou et al. 2022b; Wang et al. 2021a) across various applications. Research has expanded to open-world and open-vocabulary segmentation (Yuan et al. 2024a; Qi et al. 2022a, b), as well as unified architectures for multiple tasks (Athar et al. 2023; Jain et al. 2023; Li et al. 2024d). However, most methods focus solely on visual segmentation and lack interactive textual and visual prompts found in MLLMs. To address this, we combine SAM with MLLMs, extending SAM from segment anything to any segmentation, and introduce a unified framework adaptable to all image segmentation tasks, establishing a strong baseline.
Method
Figure 3: The Overview of X-SAM. X-SAM comprises dual encoders, dual projectors, a language model, a segmentation connector, and a segmentation decoder. The dual encoders process the image and project features to match text embedding dimensions, which are then input to the language model with tokenized text for instruction-guided understanding. The SAM features are connected to the segmentation decoder, which uses the LLM’s <SEG> token to generate segmentation masks.
@@
Vision Query Input. Beyond text query inputs, some tasks necessitate vision query inputs, such as interactive segmentation (Zhang et al. 2024c) and the Visual GrounDed segmentation proposed in this work. In contrast to text query inputs, vision query inputs incorporate a visual prompt from the user, which may take the form of points, scribbles, boxes, or masks. To denote the visual prompt, we employ a dedicated token, <region>, within the input format. Analogous to the text query input, the visual prompt is formatted as “<p><region></p>” and the segmentation output is similarly indicated by the <SEG> token. The <region> token serves as a placeholder for the visual prompt and will be replaced by the region feature extracted from the segmentation encoder.
@@
Unified Formulation. The latent language embeddings between the <p> and </p> tokens are used as the condition embeddings for the segmentation decoder to compute the classification scores. Based on this formulation, we achieve a unified framework for all image segmentation tasks. Given an input image $\mathbf{X}_{\mathrm{v}}\in\mathbb{R}^{H\times W\times 3}$ and a language instruction $\mathbf{X}_{\mathrm{q}}\in\mathbb{R}^{P\times 1}$, the model takes the image and language instruction as inputs and outputs a language response $\mathbf{Y}_{\mathrm{q}}\in\mathbb{R}^{L\times 1}$ and a segmentation mask $\mathbf{Y}_{\mathrm{m}}\in\mathbb{R}^{H\times W}$. Here, $P$ is the length of the input text tokens, and $L$ is the total length of the input and output text tokens. $H$ and $W$ denote the height and width of the image, respectively. Detailed input format examples can be found in Figure 2 (a) and (b).
Figure 4: The Architecture of Segmentation Connector.
Architecture
@@
In this section, we propose X-SAM, a unified segmentation MLLM for any segmentation. As shown in Figure 3, it includes dual encoders, dual projectors, an LLM, a segmentation connector, and a segmentation decoder.
@@
Table 2: Comprehensive Performance Comparison. We compare X-SAM to segmentation-specific models (Gray) and MLLMs. “✗” denotes unsupported tasks. “–” indicates unreported results. X-SAM achieves state-of-the-art performance across all segmentation tasks with a single model. Best results are in bold, second-best are underlined.
Dual Projectors. To enhance the LLM’s understanding of the image, we concatenate the features from the image encoder and the segmentation encoder before passing them to the LLM. Specifically, the feature from the segmentation encoder is too large to be processed directly by the LLM, so we utilize a pixel-shuffle operation to reduce its spatial size. We then project the reduced feature into the language embedding space $\mathbf{H}_{\textit{q}}$ via an MLP projector $\mathbf{\textit{W}}_{\textit{s}}$. For the feature from the image encoder, we directly project it into the language embedding space via an MLP projector $\mathbf{\textit{W}}_{\textit{i}}$, such that $\mathbf{H}_{\textit{v}}=\mathbf{\textit{W}}_{\textit{i}}\cdot\mathbf{Z}_{\texttt{v}}$ and $\mathbf{H}_{\textit{s}}=\mathbf{\textit{W}}_{\textit{s}}\cdot\mathbf{Z}_{\texttt{s}}$. We then concatenate the features from dual projectors and the language embeddings, and input them into the LLM $f_{\phi}$.
Segmentation Connector. For image segmentation tasks, fine-grained multi-scale features are crucial for the segmentation decoder to accurately predict segmentation masks. The output of the segmentation encoder in SAM is single-scale (1/16) with reduced spatial resolution. To obtain multi-scale features, we design a segmentation connector $g_{\texttt{c}}$, to bridge the segmentation encoder and decoder. As shown in Figure 4, we perform patch-merge using a pixel-shuffle (Chen et al. 2024) with a scale of 0.5 to reduce the spatial size of the last feature in the encoder to a smaller scale (1/32). We also perform patch-expand with a pixel-shuffle of scale 2.0 to increase the spatial size of the last feature to a larger scale (1/8), resulting in multi-scale features for the segmentation decoder.
Segmentation Decoder. The Segment Anything Model (SAM) can segment a single object based on input text or visual prompts, but it fails to segment all objects in a single inference. To segment all objects at once, we replace its original segmentation decoder with a new decoder, following the approach in (Cheng et al. 2022a; VS et al. 2024). The segmentation decoder $g_{\psi}$ predicts masks and their category probabilities from either the input latent embedding $\mathbf{E}_{\mathrm{i}}$ or the output latent embedding $\mathbf{E}_{\mathrm{o}}$, multi-scale segmentation features $\mathbf{F}_{\mathrm{c}}$, and a set of mask query tokens plus the <SEG> token embedding, which bridges the LLM output with the segmentation decoder. Notably, we introduce a latent background embedding to represent the “ignore” category for all tasks, thereby unifying all image segmentation tasks with one model.
Method (M)LLM RefCOCO RefCOCO+ RefCOCOg
val testA testB val testA testB val test
SEEM-L (Zou et al. 2023b) - - - - - - - 65.6 -
UNINEXT-L (Yan et al. 2023) - 80.3 82.6 77.8 70.0 74.9 62.6 73.4 73.7
UNINEXT-H (Yan et al. 2023) - 82.2 83.4 81.3 72.5 76.4 66.2 74.7 76.4
OMG-LLaVA (Zhang et al. 2024b) InternLM-7B 77.2 79.8 74.1 68.7 73.0 61.6 71.7 71.9
Sa2VA(Yuan et al. 2025) InternVL2-8B 81.6 - - 76.2 - - 78.7 -
PSALM (Zhang et al. 2024c) Phi-1.5-1.3B 83.6 84.7 81.6 72.9 75.5 70.1 73.8 74.4
HyperSeg (Wei et al. 2024b) Mipha-3B 84.8 85.7 83.4 79.0 83.5 75.2 79.4 78.9
X-SAM (Ours) Phi-3-3.8B 85.1 87.1 83.4 78.0 81.0 74.4 83.8 83.9
Table 3: Comparison of Referring Segmentation. We evaluate methods on referring segmentation benchmarks by (M)LLMs.
Training
@@
Stage 2: Alignment Pre-training. To align the language embeddings and visual embeddings, we perform alignment pre-training on the LLaVA-558K dataset, following (Liu et al. 2023a). We keep the dual encoders and the LLM parameters frozen, and only train the dual projectors. In this way, the image embeddings and segmentation embeddings can be aligned with the pre-trained LLM word embeddings. The training objective for alignment pre-training is an auto-regressive loss $\mathcal{L}_{\mathrm{regressive}}$:
@@
Experiments
Experiment Settings
Datasets and Tasks. For segmentor fine-tuning, we train on the COCO-Panoptic (Kirillov et al. 2019) dataset. For alignment pre-training, we utilize the LLaVA-558K (Liu et al. 2023a) dataset. For end-to-end mixed fine-tuning, we incorporate one image conversation dataset and five types of image segmentation datasets into the training process. To balance the training data across these diverse datasets, we set the training epoch to 1 and adjust the resampling rates of different datasets using dataset balance resampling. After training, X-SAM is capable of performing a variety of tasks, including Image Conversation, Generic, Referring, Reasoning, GCG, Interactive, and VGD Segmentation. Additionally, X-SAM supports Open-Vocabulary (OV) (OV-semantic, OV-instance, OV-panoptic) segmentation, enabling it to segment all objects defined by the input prompt, even those never seen before. Note that COCO-VGD is our proposed VGD segmentation dataset, which is built on the COCO2017 dataset.
Evaluation Metrics. We conduct extensive experiments to evaluate the performance of X-SAM. For generic segmentation and open-vocabulary segmentation, we use PQ, mIoU, and mAP as the main metrics for panoptic, semantic, and instance segmentation, respectively. For referring segmentation and reasoning segmentation, we adopt cIoU and gIoU as metrics, following (Zhang et al. 2024c). For GCG segmentation, we use M, C, AP50, and mIoU as metrics, following (Rasheed et al. 2024). For interactive segmentation, we use mIoU and cIoU, also following (Zhang et al. 2024c). For VGD segmentation, we use AP and AP50. For image conversation, we adopt scores from common MLLM benchmarks as the main metrics, following (Liu et al. 2023a).
Implementation Details. We adopt the XTuner (Contributors 2023) codebase for training and evaluation. During segmentor fine-tuning, we train all parameters, set the batch size to 64, and use a learning rate of 1e-5 for the SAM encoder and 1e-4 for the other parameters. The number of training epochs is set to 36. For alignment pre-training, we train only the dual projector parameters, with a batch size of 256, a learning rate of 1e-3, and one training epoch. For end-to-end mixed fine-tuning, we train all parameters, set the batch size to 64, and use a learning rate of 4e-6 for the dual encoders and 4e-5 for the other parameters, with one training epoch. All training is conducted on 16 A100 GPUs. For image conversation evaluation, we use the VLMEvalKit (Duan et al. 2024) codebase to evaluate performance on MLLM benchmarks. For segmentation task evaluation, we follow the settings described in the corresponding papers and repositories.
@@
Table 4: Comparison of GCG Segmentation. $\dagger$ indicates pretraining with the GranD dataset (Rasheed et al. 2024).
Main Results
@@
Table 5: Comparison of VGD Segmentation. $\dagger$ indicates evaluation results following X-SAM setting.
@@
Table 6: Ablation on Fine-Tuning(FT).
@@
Table 7: Ablation on Dual Encoders. Swin is initialized from Mask2Former (M2F) (Cheng et al. 2022a).
@@
Overall. In Table 2, we compare X-SAM with current segmentation-specific models and MLLMs. X-SAM demonstrates the most comprehensive capabilities. It achieves performance comparable to state-of-the-art in generic segmentation, and achieves the best performance on other benchmarks, with a single model. X-SAM sets a new state-of-the-art record for image segmentation benchmarks. Detailed results for each task are discussed below.
Referring Segmentation. We evaluate X-SAM on RefCOCO, RefCOCO+, and RefCOCOg, with the results shown in Table 3. X-SAM outperforms PSALM (Zhang et al. 2024c) by 1.5% cIoU, 5.1% cIoU, and 10.0% cIoU on the validation sets of RefCOCO, RefCOCO+, and RefCOCOg, respectively. Compared to Sa2VA-8B (Yuan et al. 2025), X-SAM achieves better results with a smaller model size. It shows performance improvements of 3.5% cIoU, 1.8% cIoU, and 5.1% cIoU on RefCOCO, RefCOCO+, and RefCOCOg, respectively.
GCG Segmentation. Grounded conversation generation demands detailed image and pixel-level understanding, requiring MLLMs to link captioned objects to their segmentation masks. As shown in Table 4, X-SAM achieves a significant performance improvement compared to previous methods and obtains the best results on both the Val and Test sets. In terms of image-level understanding, X-SAM outperforms GLaMM (Rasheed et al. 2024) by 0.2% METEOR and 3.2% CIDEr on the Val set, and by 0.5% METEOR and 4.8% CIDEr on the Test set. In terms of pixel-level understanding, X-SAM outperforms OMG-LLaVA (Zhang et al. 2024b) by 3.3% AP and 3.9% mIoU on the Val set, and by 4.3% AP and 4.3% mIoU on the Test set.
VGD Segmentation. Visual grounded segmentation demands vision query understanding, requiring MLLMs to comprehend the visual modality and segment all related instances. Table 5 presents the VGD segmentation results. As VGD segmentation is our newly proposed task, we evaluate PSALM (Zhang et al. 2024c) following X-SAM’s settings. X-SAM outperforms PSALM by 45.9% AP, 45.9% AP, 45.8% AP, and 47.4% AP on Point, Scribble, Box, and Mask visual prompts, respectively.
Abaltions